From: STOSC::SMM "On-Duty SMM x5074" 22-JAN-1993 21:59:43.46 To: @[.DIS]HST_STAT CC: Subj: HST STATUS REPORT, 01/21 - 01/22. HUBBLE SPACE TELESCOPE DAILY REPORT #807 PERIOD COVERED: 0800 EST 01/21/93 - 0800 EST 01/22/93 Daily Status Report as of 022/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed FOS/RD/BL 4694 (Stellar Content of Galaxies and Globular Clusters) The Faint Object Spectrograph was used to make three observations of four globular clusters located in the central bulge of galaxy M-31 (aka, Andromeda Galaxy). The observations were received in real time. The peakups were successful, and the science observations looked good. 1.2 Completed WFC 3917 (HST Medium Deep Survey: Part 1) The Wide Field Camera was used to make observations for a high latitude medium-deep sky survey. The observations were taken in parallel with the prime science observations. The proposal was executed as scheduled, and no problems were reported. 1.3 Completed Two Sets of FOS 4696 (Wave Length Calibration: Internal /External Offsets) The Faint Object Spectrograph was used to take a series wave length calibration observations using both external and internal sources. This is to determine the offsets between the respective wave length scales. The observations were received and displayed for the principal investigator, who was present in the OSS. All observations looked normal. 1.4 Completed FGS 4031 (Search for Planetary Companions to Low Mass) The Fine Guidance Sensors were used to measure the position of the star Proxima Centauri relative to background field stars. The intent is to determine the presence of a Jupiter sized planet in orbit around the target star. The observations were executed as planned, and no problems were reported. 1.5 Completed FOS/BL 4045 (UV Spectroscopy of Low Redshift Galaxies) The Faint Object Spectrograph was used to make ultraviolet spectral observations of the active galactic nucleus of galaxy FAIRALL9. The principal investigator (PI) was present to review the data. The PI indicated that the data was very good. 1.6 Uplinked Updated Ephemris Tables, Per MOSES 2871 1.7 Completed PC 4140 (Rapid Internal Monitor, Part 2) The Planetary Camera was used to make internal flat observations to check for new contamination and the reappearance of quantum efficiency hysteresis (QEH). The observations were executed as scheduled, and no problems were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 4 Successful : 4 Scheduled Reacquisitions : 12 Successful : 12 Losses of Lock : 0 2.2 FHST Updates: Scheduled : 9 Successful : 9 2.3 Significant HSTARs: None 2.4 Operations Note: None 3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD: 3.1 Complete GHRS/UV2 1095 (Variability of High Luminosity Stars) 3.2 Complete GHRS/2 3578 (Line Eclipse Mapping of Accretion Disk Wind) 3.3 Complete GHRS/2 4462 (SSA A2 Offset Verification) 3.4 Complete PC 4347 (Flood Retention Test) 3.5 Complete Two Sets of GHRS/2 2537 (Physical Conditions of Interstellar Gas) 3.6 Uplink Updated Ephemeris Tables, Per MOSES 2871 3.7 Complete WFC-PC 4109 (Rapid Internal Monitor) 3.8 Complete Two Sets of PC 3848 (Separate Light Curves of Pluto and Charon) 3.9 Complete FOS/RD 4052 (Inner Regions of Quasars) 3.10 Complete WFC 3917 (HST Medium Deep Survey, Part 1) 3.11 Complete WFC 4216 (Calibration of Supernova) 3.12 Complete WFC 2227 (Determine Extragalactic Distance Scale: M-81) 3.13 Complete GHRS/2 3444 (Gas in Individual Interstellar Clouds) 4.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: None 5.0 SIGNIFICANT FORTHCOMING EVENTS: The Zero-Gyro (On-Orbit) Test will be held from Tuesday 026/1200Z through Thursday 028/0000Z. The Fight Readiness Review meeting (FRR) was held today Friday Jan. 22 .